Assume the following production function: Y=8K^(0.3)N^(0.7)

a) Keeping capital constant, show graphically a technological improvement. (10 marks)

b) If the capital stock is K = 30, and the prevailing wage rate is w* = $25, find the full employment. (10 marks)
Answer #1

We have a production function: Y = 8K^(0.3)N^(0.7) where technology parameter is 8.

a) Keeping capital constant, If there is a technological improvement, it will increase the overall productivity of both labor and capital. Now that capital is fixed, and only labor is variable, with increase in labor units, the new production function will lie above the original one, assuming that technology parameter is increased to 12 and production function is Y = 12K^(0.3)N^(0.7)

b) If the capital stock is K = 30, production function Y = 8*(30^(0.3)*(N^(0.7) = 22.1935(N^(0.7) and the prevailing wage rate is w* = $25, find the MPN

MPN = dY/dN = 0.7*22.1935/N^0.3

At the optimal allocation, MPN = wage rate

25 = 15.5345/N^0.3 and N = 0.204.

This is the optimum amount of labor

Total Output Y A =8 LI Labor Force, L